# Interpretable Machine Learning Models for Predicting the Next Targets of Activist Funds

### v1

@@ Abstract
activist funds , shareholder activism , corporate finance , machine learning , SHAP
@@ 1 Introduction
Activist funds are a distinct class of investment funds that seek to effect change within their portfolio companies, not just through investment but by actively influencing management and strategic decisions. These funds differentiate themselves from traditional investment vehicles by adopting a hands-on approach, leveraging their stakes in companies to advocate for changes in corporate governance, strategy, management, and operations ([2]). Their ultimate goal is to unlock shareholder value that might not be realized under the current management’s approach.
@@
The participation of activists has been growing recent years, particularly after the financial scandals and the financial crisis of 2007-2008 ([12]). In 2018, nearly one campaign launched against a new target every day ([19]). Thanks to regulatory reforms that have made shareholder engagement more accessible and weakened corporate defenses, activists can now more easily challenge management to enhance shareholder benefits ([9]). The motivations behind activist investing are diverse. Romito identifies four main types of activist investors ([25]). Some focus on enhancing returns by advising management to issue debt or redirect free cash flow toward dividends or stock buybacks. Others advocate for merging similar competitors to capitalize on economies of scale. Another group prefers to enhance profitability by separating mismatched divisions within a company. Additionally, some investors aim to streamline operations by maintaining only the most profitable units and spinning off less effective ones. Beyond these financial goals, certain socially responsible funds also promote environmental, social, and governance (ESG) improvements, aiming to foster ethical business practices ([4]).
@@
Studies have demonstrated the benefits of activist interventions in terms of market value, profitability, and governance, although there is ongoing debate about how these benefits vary over time ([12], [14], [3]). Bebchuk et al. conducted an analysis of approximately 2,000 activist interventions from 1994 to 2007, revealing that activist funds often target underperforming companies with stock returns significantly lower than those of industry peers. However, in the five years following these interventions, the performance of the targeted companies generally improved, with them closing two-thirds of their performance gap in terms of return on assets compared to their peers ([3]). Conversely, Desjardine and Durand analyzed 1,324 campaigns between 2000 and 2016 and found that while shareholder benefits were initially apparent and significant, they were short-lived. These gains often came at a long-term cost to other stakeholders, manifested in reduced cash flow and lower investment spending ([12]).
@@
Recognizing potential targets of these activist funds is essential for a wide array of stakeholders. It enables corporations to prepare and defend against unwanted interventions, guides activist investors in selecting firms where they can significantly impact, and offers retail investors opportunities for profit through potential stock price appreciations following activist campaigns.
@@
This study is focused on creating a model that can accurately predict which companies will become targets of activist funds. We analyzed data from firms listed on the Russell 3000 index from 2016 to 2022, along with information on activist campaigns. Our analysis included a wide range of variables encompassing traditional measures of valuation and operations, as well as alternative data points related to ownership structures, governance, and technical indicators. By experimenting with 123 distinct combinations of imputation, oversampling, and machine learning techniques, we were able to determine the most effective model. Furthermore, we utilized Shapley values ([20]) to pinpoint the key factors that influence a company’s likelihood of being targeted by an activist fund. This approach not only enhances the predictive accuracy of our model but also provides deeper insights into the dynamics of shareholder activism.
@@ 2 Literature Review
Several studies explored the valuation analysis strategies of activist funds. Pfirrmann and Eichner’s study based on 30 campaigns between 2013 and 2019, reveals a preference for relative valuation over intrinsic valuations, primarily utilizing primarily utilizing Enterprise Value/EBITDA (EV/EBITDA) and Price/Earnings (P/E) ratios ([22]). Consistent with findings by Brav et al. ([5]) and Boyson and Mooradian ([7]), it is observed that firms targeted by activists often exhibit lower valuations compared to their fundamental values. Operationally, these firms show robust cash flows from operations and sales growth rates. Furthermore, their asset returns outperform those of their competitors, despite their stock performance trailing behind the broader market. The analysis also highlights that activist investors typically secure a significant stake in these firms, purchasing between 5% to 10% of equity shares on the market once they have identified their targets.
@@
We broadened the scope of our investigation to how general institutional investors approach company valuation. Bancel and Mittoo conducted a survey among 365 European finance experts holding CFAs or similar qualifications, finding that the most favored valuation techniques are Discounted Cash Flows (DCF) and relative valuation methods, particularly emphasizing the EV/EBITDA and P/E ratios as the top choices for multiples ([8]). Similarly, Mukhlynina and Nyborg’s (2016) study involving valuation specialists, such as investment bankers, consultants, and private equity professionals, revealed that a significant portion of these experts employ both multiples and DCF in their valuation practices, with 47% admitting to using both methods but predominantly favoring multiples ([21]).
@@
Several studies have explored the evolving focus of activist funds on environmental, social, and governance (ESG) issues. A study by Zhu revealed that activist interventions not only improve the valuations and operations of a company but also steer management practices in a more shareholder-friendly direction. These interventions have led to increased shareholder returns through higher buyback activities and increased dividend payout ratios, as well as reductions in CEO compensation ([29]). In addtion, Barko et al. demonstrated that socially responsible activist funds often choose large, high-profile companies with good financial performance and liquidity but poor ESG scores. Their research indicates that activism focused on corporate social responsibility typically enhances both ESG practices and corporate sales performance, suggesting that ethical investing can coincide with robust financial returns ([4]). The trend towards social responsibility in activism is further supported by research from Albuquerque et al., who developed a model demonstrating that firms decide to engage in CSR activities. Their study noted that over 3,000 institutions, managing assets upwards of 90 trillion USD, have committed to integrating CSR into their due diligence processes ([1]). Additionally, research by Francis et al. underscored the significance of governance-related criteria in target selection, revealing that activist funds are approximately 52% more likely to choose firms with female CEOs. These targets led by women are more cooperative towards activists and tend to see greater improvements in market and operational performance during campaigns ([13]).
Our research builds upon previous studies to provide fresh insights into shareholder activism. We consider various aspects of companies, including valuations, operations, governance, ownership, stock performance, and technical information, to offer a comprehensive analysis of target selection. Additionally, we employ multiple machine learning techniques to develop a model that predicts potential future targets and to identify the most critical information for making these predictions.
@@ 3.1 Data Collection
In this study, we analyze the impact of 46 independent variables related to corporate characteristics on the likelihood of a company being targeted by activist investors. These variables are categorized as follows: 10 governance indicators, 3 ownership metrics, 6 technical indicators, 7 return measures, 10 valuation metrics, and 10 operational indicators, with details presented in Table 1. The independent variables, sourced from the Bloomberg Terminal, comprises year-end snapshots for companies listed on the Russell 3000 index from 2016 to 2022. For the dependent variable that indicates if a company is targeted or not, we utilize data on activist campaigns reported from 2011 to 2022, also obtained from the Bloomberg Terminal. A company is considered a target if it becomes the subject of an activist campaign within the subsequent 12 months. Firms actively under a campaign are temporarily excluded from the dataset and reinstated after the campaign concludes. Our final dataset consists of 19,414 instances, with approximately 3.4% of these companies being identified as targets (see Fig. 1).
Figure 1: From 2016 to 2022, around 2 to 4% of the companies in Russell 3000 were targeted by activist funds, with the mean value of 3.4% in the entire period.
@@ 3.2.2 Data Imputation
There is a significant portion of missing values in the data set (see Fig. 2). For data imputation, we use 4 different methods: mean imputation, K-nearest neighbours (KNN) imputations (k=5) ([6]), Multivariate Imputation by Chained Equations using LightGBM (MiceForest)([26]), and Generative Adversarial Imputation Networks (GAIN) ([27]). K-nearest neighbors (KNN) imputation is a technique used to fill in missing values in a dataset based on the values of its neighboring data points. The missing value is estimated by averaging or taking a weighted average of the values of the k nearest neighbors to the missing point. Multivariate Imputation by Chained Equations (MICE) is a method used to impute missing values in a dataset by iteratively modeling each variable with missing data as a function of other variables in the dataset. Using LightGBM (Light Gradient Boosting Machine) for MICE involves employing gradient boosting, a powerful machine learning technique, to predict missing values iteratively. LightGBM is particularly efficient for this task due to its ability to handle large datasets and its fast training speed. Generative Adversarial Imputation Networks (GAIN) is a technique used for imputing missing values in a dataset using deep learning and generative adversarial networks (GAN). GAIN is particularly effective for imputing missing values in high-dimensional datasets with complex dependencies between variables.
It’s crucial to highlight that for all imputation methods except for mean imputation, we perform the imputation process within the same category of variables, incorporating the one-hot encoded values of the companies’ year. This approach ensures that imputations are conducted among variables that share predictive relationships and also accounts for overarching yearly trends. To avoid target leakage, we apply these data imputation techniques exclusively to the training set. For the test sets, we insist on median imputation.
@@
Board size Large boards may signal poor governance or conflict of interest.
@@
30-day Relative Strength Index Low RSI indicates stock is oversold; provides favorable entry
@@
3.2.3 Data Oversampling
In addressing the challenge posed by the dataset, where only 3.41% of the entries are labeled as targeted, we confronted a significant imbalance between the majority and minority classes. To mitigate this issue and enhance the robustness of our analysis, we implement 4 oversampling techniques. These include Random Oversampling Example (ROSE) ([28]), Synthetic Minority Oversampling Technique (SMOTE) ([10]), borderline SMOTE, and Adaptive Synthetic Sampling (ADASYN) ([15]). Random oversampling (ROSE) performs a random duplicating instances from the minority class(es) until the class distribution is more balanced. SMOTE is a popular method used to address class imbalance in datasets, particularly in the context of binary classification problems. Unlike random oversampling, which simply duplicates instances from the minority class, SMOTE generates synthetic samples for the minority class by interpolating between existing minority class instances. By creating synthetic samples instead of simply duplicating existing minority class instances, SMOTE helps to increase the diversity of the minority class and mitigate the risk of overfitting.Borderline SMOTE and ADASYN are two extensions of SMOTE designed to further address class imbalance in datasets. While SMOTE generates synthetic samples uniformly across the minority class, Borderline SMOTE focuses specifically on those instances that are close to the decision boundary — often the most difficult to classify. It primarily generates synthetic examples near these borderline cases, aiming to strengthen the decision boundary for the classifier. On the other hand, ADASYN adjusts the number of synthetic samples generated for each minority class instance according to their level of difficulty in learning, where more samples are generated for instances harder to learn. This adaptive approach helps in focusing on the regions where the classifier is most likely to benefit from more diverse data.
In our case, oversampling is only applied to training set in order to maintain purity of testing set.
Figure 2: The first subfigure shows the number of instances with different number of missing values for 46 variables. The second subfigure shows the proportion of missing values for each variable.
@@ 3.3 Model Training & Metrics
We employ six diverse machine learning models in our analysis: logistic regression ([18]), random forest ([24]), XGBoost ([11]), LightGBM ([16]), CatBoost ([23]), and neural networks ([17]). The data is divided, allocating 80% (15,531 instances) for training—this figure increases when oversampling is applied—and the remaining 20% (3,883 instances) for testing. By integrating 4 imputation methods, 5 oversampling strategies (including the option of not oversampling), and 6 machine learning models, we arrive at 120 distinct configurations. Additionally, recognizing the capabilities of LightGBM, XGBoost, and CatBoost to efficiently handle sparse data, we account for these as three extra models using original data without imputation, increasing our total to 123 models.
@@ 3.4 Model Interpretation Method
In our study, we employed SHAP (SHapley Additive exPlanations) to analyze and interpret the contributions of different features in our models. Grounded in cooperative game theory, SHAP provided a robust framework for explaining the output of any machine learning model ([20]). We chose this method primarily for two reasons. First, the effectiveness of SHAP in highlighting the relative importance of variables was extensively validated in previous research. Second, unlike model-specific tools such as coefficients in linear regression or MDI (mean decrease in impurity) and the number of splits in tree-based models like Random Forest or LightGBM, SHAP is model-agnostic. This universality made it a versatile tool for application across various model architectures.
@@ 4.1 Model Performance
In an evaluation of 123 different combinations of imputation, oversampling, and machine learning methods, 11 achieved an AUC-ROC score above 0.7, with 4 surpassing 0.75 in the test set (see table 2). The highest AUC-ROC score observed was 0.782, achieved with a strategy combining KNN imputation, Borderline SMOTE oversampling, and logistic regression. This study highlights the superior performance of machine-learning-based imputation techniques over simpler strategies like median imputation, which was used in only one of the top 11 combinations. Borderline SMOTE was particularly effective, being implemented in 10 of the 11 highest-scoring models. Logistic regression emerged as the preferred method in all four top-performing combinations with AUC-ROC scores above 0.75, whereas ensemble methods were used in the remaining seven models with scores below 0.75. No neural network models scored above 0.7, suggesting that logistic regression may offer more robustness in smaller datasets compared to more complex methods like ensemble approaches and neural networks.
Table 2: Summary of $1$ models that have AUC-ROC (test set) of greater than 0.7

### v2

# ORCID: 0009-0005-8104-2142

@@ Abstract
Activist Funds , Shareholder Activism , Corporate Finance , Machine Learning , Explainable AI , SHAP
@@ 1 Introduction
Activist funds are a distinct class of investment funds that seek to effect change within their portfolio companies, not just through investment but by actively influencing management and strategic decisions. These funds differentiate themselves from traditional investment vehicles by adopting a hands-on approach, leveraging their stakes in companies to advocate for changes in corporate governance, strategy, management, and operations [2]. Their ultimate goal is to unlock shareholder value that might not be realized under the current management’s approach.
@@
The participation of activists has been growing recent years, particularly after the financial scandals and the financial crisis of 2007-2008 [12]. In 2018, nearly one campaign launched against a new target every day [19]. Thanks to regulatory reforms that have made shareholder engagement more accessible and weakened corporate defenses, activists can now more easily challenge management to enhance shareholder benefits [9]. The motivations behind activist investing are diverse. Romito identifies four main types of activist investors [27]. Some focus on enhancing returns by advising management to issue debt or redirect free cash flow toward dividends or stock buybacks. Others advocate for merging similar competitors to capitalize on economies of scale. Another group prefers to enhance profitability by separating mismatched divisions within a company. Additionally, some investors aim to streamline operations by maintaining only the most profitable units and spinning off less effective ones. Beyond these financial goals, certain socially responsible funds also promote environmental, social, and governance (ESG) improvements, aiming to foster ethical business practices [4].
@@
Studies have demonstrated the benefits of activist interventions in terms of market value, profitability, and governance, although there is ongoing debate about how these benefits vary over time [12], [13], [6]. Bebchuk et al. conducted an analysis of approximately 2,000 activist interventions from 1994 to 2007, revealing that activist funds often target underperforming companies with stock returns significantly lower than those of industry peers. However, in the five years following these interventions, the performance of the targeted companies generally improved, with them closing two-thirds of their performance gap in terms of return on assets compared to their peers [6]. Conversely, Desjardine and Durand analyzed 1,324 campaigns between 2000 and 2016 and found that while shareholder benefits were initially apparent and significant, they were short-lived. These gains often came at a long-term cost to other stakeholders, manifested in reduced cash flow and lower investment spending [12].
@@
Recognizing potential targets of these activist funds is therefore crucial for a wide array of stakeholders in today’s market regime. It enables corporations to prepare and defend against unwanted interventions, guides activist investors in selecting firms where they can significantly impact, and offers retail investors opportunities for profit through potential stock price appreciations following activist campaigns.
@@
Our research focuses on creating a model that can accurately predict which companies will become targets of activist funds. We analyze data from firms listed on the Russell 3000 index from 2016 to 2022, along with information on activist campaigns. Our analysis includes a wide range of variables encompassing traditional measures of valuation and operations, as well as alternative data points related to ownership structures, governance, and technical indicators. By experimenting with 123 distinct combinations of imputation, oversampling, and machine learning techniques, we determine the most effective model. Furthermore, we utilize Shapley values [21] to pinpoint the key factors that influence a company’s likelihood of being targeted by an activist fund. This approach not only enhances the predictive accuracy of our model but also provides deeper insights into the dynamics of shareholder activism.
@@ 2 Literature Review
Activist funds have long been known to focus on traditional valuation metrics when selecting targets, as revealed by various studies. Pfirrmann and Eichner’s study based on 30 campaigns between 2013 and 2019, reveals a preference for relative valuation over intrinsic valuations, primarily utilizing Enterprise Value/EBITDA (EV/EBITDA) and Price/Earnings (P/E) ratios [24]. Consistent with findings by Brav et al. [8] and Boyson and Mooradian [7], it is observed that firms targeted by activists often exhibit lower valuations compared to their fundamental values. Operationally, these firms show robust cash flows from operations and sales growth rates. Furthermore, their asset returns outperform those of their competitors, despite their stock performance trailing behind the broader market. The analysis also highlights that activist investors typically secure a significant stake in these firms, purchasing between 5
@@
In order to reconfirm the preference for traditional valuation metrics, the evaluation methods used by institutional investors, not just activist funds, were also examined, and similar trends were confirmed in various studies. Bancel and Mittoo conducted a survey among 365 European finance experts holding CFAs or similar qualifications, finding that the most favored valuation techniques are Discounted Cash Flows (DCF) and relative valuation methods, particularly emphasizing the EV/EBITDA and P/E ratios as the top choices for multiples [3]. Similarly, Mukhlynina and Nyborg’s (2016) study involving valuation specialists, such as investment bankers, consultants, and private equity professionals, revealed that a significant portion of these experts employ both multiples and DCF in their valuation practices, with 47% admitting to using both methods but predominantly favoring multiples [23].
@@
Several studies have explored the evolving focus of activist funds on environmental, social, and governance (ESG) issues. A study by Zhu revealed that activist interventions not only improve the valuations and operations of a company but also steer management practices in a more shareholder-friendly direction. These interventions have led to increased shareholder returns through higher buyback activities and increased dividend payout ratios, as well as reductions in CEO compensation [33]. In addtion, Barko et al. demonstrated that socially responsible activist funds often choose large, high-profile companies with good financial performance and liquidity but poor ESG scores. Their research indicates that activism focused on corporate social responsibility typically enhances both ESG practices and corporate sales performance, suggesting that ethical investing can coincide with robust financial returns [4]. The trend towards social responsibility in activism is further supported by research from Albuquerque et al., who developed a model demonstrating that firms decide to engage in CSR activities. Their study noted that over 3,000 institutions, managing assets upwards of 90 trillion USD, have committed to integrating CSR into their due diligence processes [1]. Additionally, research by Francis et al. underscored the significance of governance-related criteria in target selection, revealing that activist funds are approximately 52% more likely to choose firms with female CEOs. These targets led by women are more cooperative towards activists and tend to see greater improvements in market and operational performance during campaigns [14].
Our research builds upon previous studies to provide fresh insights into shareholder activism. Specifically, the research have 4 main contributions:
  • 1.
    First, we explore a wide range of company characteristics to assess their likelihood of being targeted by activist funds. While traditional variables like valuations and operations have been extensively studied, we incorporate less-explored factors such as governance, ownership, technical, and social aspects. This offers a more comprehensive view of the companies that attract activist interest.
  • 2.
    Second, we employ advanced machine learning techniques to develop a predictive model. Unlike traditional econometric methods like logistic regression, which are limited in capturing the complexity of activist decision-making, machine learning models can exploit non-linear and non-parametric patterns, resulting in significantly improved prediction accuracy.
  • 3.
    Third, we go beyond conventional methods of analyzing variable importance, like standard coefficient analysis, by using SHAP (SHapley Additive exPlanations). This game-theory-based approach allows us to measure the contribution of different variables at the individual instance level, providing deeper insights into how activists select their targets.
  • 4.
    Fourth, we analyze the most recent data from the US equity market, covering the period from 2016 to 2022. Given the dynamic nature of financial markets and frequent regime changes, focusing on this latest data enables us to capture the most relevant and up-to-date trends in activist fund behavior.
@@ 3 Methodology
Figure 1 provides an overview of the proposed pipeline for activist investment target prediction, consisting of several stages. It begins by collecting data from Russell 3000 companies (2016-2022) and activist campaigns (2011-2023), featuring diverse variables. In preprocessing, non-ratio variables are transformed into ratios and selected variables are converted to percentiles relative to industry peers. Four imputation techniques handle missing data, and four oversampling methods address class imbalance. Six machine learning models are trained using these imputed and balanced datasets. SHAP values are then used to interpret the models, highlighting key variables. Each step of this framework is detailed in the following subsections.
Figure 1: Overview of the methodology.
@@ 3.1 Data Collection
In this study, we analyze the impact of 46 independent variables related to corporate characteristics on the likelihood of a company being targeted by activist investors. These variables are categorized as follows: 10 governance indicators, 3 ownership metrics, 6 technical indicators, 7 return measures, 10 valuation metrics, and 10 operational indicators, with details presented in Table 1.
The independent variables, sourced from the Bloomberg Terminal, comprises year-end snapshots for companies listed on the Russell 3000 index from 2016 to 2022. We chose the Russell 3000 index as the subject of this study because it represents approximately 98% of the investable U.S. equity market, covering both large-cap and small- to mid-cap companies  [^1] https://research.ftserussell.com/Analytics/FactSheets/Home/DownloadSingleIssue?issueName=US3000USD&isManual=True. Additionally, the U.S. market provides the most comprehensive data on corporate characteristics, and is the market where activist campaigns are the most active, making it an ideal setting for our analysis. For the dependent variable that indicates if a company is targeted or not, we utilize data on activist campaigns reported from 2011 to 2023, also obtained from the Bloomberg Terminal. A company is considered a target if it becomes the subject of an activist campaign within the subsequent 12 months. Firms actively under a campaign are temporarily excluded from the dataset and reinstated after the campaign concludes. Our final dataset consists of 19,414 instances, with approximately 3.4% of these companies being identified as targets (see Fig. 2).
Figure 2: Distributions of the targeted and non-targeted companies in the dataset. From 2016 to 2022, around 2% to 4% of the companies in Russell 3000 were targeted by activist funds, with the mean value of 3.4% in the entire period.
@@ 3.2.2 Data Imputation
Missing data is a common issue in data science, arising from factors like human error or incomplete records, and it can impede analysis. Imputation methods provide a systematic approach to estimate and fill in these gaps, allowing for more robust data analysis without the need to discard incomplete observations. We use four different methods: Mean Imputation, K-nearest neighbours (KNN) imputations (k=5) [5], Multivariate Imputation by Chained Equations using LightGBM (MiceForest) [29], and Generative Adversarial Imputation Networks (GAIN) [31]. For further details of the methods, refer to Appendix B.
For all imputation methods except for mean imputation, we perform the imputation process within the same category of variables, incorporating the one-hot encoded values of the companies’ year. This approach ensures that imputations are conducted among variables that share predictive relationships and also accounts for overarching yearly trends. To prevent target leakage, we apply these data imputation techniques only to the training set. For the test set, we exclusively use median imputation.
3.2.3 Data Oversampling
In addressing the challenge posed by the dataset, where only 3.4% of the entries are labeled as targeted, we confronted a significant imbalance between the majority and minority classes. Highly imbalanced data presents an additional challenge, as most models tend to favor the majority class, and in extreme situations, they may completely overlook the minority class. To mitigate this issue and enhance the robustness of our analysis, we conduct oversampling to adjust the ratio of targeted and non-targeted entries to 50:50.
Specifically, we implement 4 oversampling techniques. These include Random Oversampling Example (ROSE) [32], Synthetic Minority Oversampling Technique (SMOTE) [10], borderline SMOTE, and Adaptive Synthetic Sampling (ADASYN) [15]. For further details of the methods, refer to Appendix C.
@@ 3.2.3 Data Oversampling
Board size Less effort needed to gain support from the smaller-sized board.
@@
30-day Relative Strength Index Low RSI indicates stock is oversold; provides favorable entry points.
@@
Figure 3: The first subfigure shows the proportion of missing values for each variable.The second subfigure shows the number of instances with different number of missing values for 46 variables.
@@ 3.3 Model Training & Metrics
We employ six diverse machine learning models in our analysis: Logistic Regression [18], Random Forest [26], XGBoost [11], LightGBM [16], CatBoost [25], and neural networks [17]. For further details of the methods, refer to Appendix D.
The data is divided, allocating 80% (15,531 instances) for training—this figure increases when oversampling is applied—and the remaining 20% (3,883 instances) for testing. By integrating 4 imputation methods, 5 oversampling strategies (including the option of not oversampling), and 6 machine learning models, we arrive at 120 distinct configurations. Additionally, recognizing the capabilities of LightGBM, XGBoost, and CatBoost to efficiently handle sparse data, we account for these as three extra models using original data without imputation, increasing our total to 123 models.
@@ 3.4 Model Interpretation Method
In our study, we employ SHAP (SHapley Additive exPlanations) to analyze and interpret the contributions of different features in our models. Grounded in cooperative game theory, SHAP provided a robust framework for explaining the output of any machine learning model [21]. We choose this method primarily for two reasons. First, the effectiveness of SHAP in highlighting the relative importance of variables was extensively validated in previous research. Second, unlike model-specific tools such as coefficients in linear regression or MDI (mean decrease in impurity) and the number of splits in tree-based models like Random Forest or LightGBM, SHAP is model-agnostic. This universality makes it a versatile tool for application across various model architectures.
@@ 4.1 Model Performance
In an evaluation of 123 different combinations of imputation, oversampling, and machine learning methods, 11 models achieved an AUC-ROC score above 0.7, with 4 surpassing 0.75 on the test set (see Table 2). The highest AUC-ROC score recorded was 0.782, attained using a combination of KNN imputation, Borderline SMOTE oversampling, and logistic regression.
Among the 11 top-performing models, 10 utilized machine learning-based imputation, with only one employing naive imputation. None of the three models capable of handling sparse data surpassed an AUC-ROC of 0.7. The result suggests that machine learning-based imputation methods significantly enhanced predictive performance.
A similar trend was observed with oversampling methods, as only one model never employed oversampling. This suggests that oversampling effectively addressed the class imbalance problem, which can often cause models to overlook or undervalue the characteristics of the minority class. Notably, all 10 oversampling techniques used in the top 11 models employed Borderline SMOTE. Borderline SMOTE targets instances near the decision boundary between the two classes. The effectiveness of Borderline SMOTE indicates the possibility of a distinct boundary between targeted and non-targeted instances in the vector space, a subject that could be further explored in future research.
In terms of machine learning methods, all but neural networks appeared among the top 11 models. All of the 4 models that surpassed an AUC-ROC score of 0.75 employed traditional logistic regression, indicating that more advanced machine learning models, such as neural networks and ensemble methods, did not lead to superior performance. This trend, where logistic regression outperforms more complex models, has been frequently observed in many studies across various domains [30], [20], [22], [28]. A possible explanation is that advanced models often require larger datasets to effectively learn the underlying patterns, while logistic regression can perform robustly even with smaller datasets. This finding emphasizes the effectiveness of logistic regression, particularly in contexts where data size is limited.
Table 2: Summary of 11 models that have AUC-ROC (test set) of greater than 0.7